# RISCORE: Enhancing In-Context Riddle Solving in Language Models through Context-Reconstructed Example Augmentation

### v3

@@ Appendix F Experimental Setup
AI@Meta (2024) In our experiments, we chose two variations of the model: the 8B and 70B versions. The Llama3-8B999meta-llama/Meta-Llama-3-8B-Instruct model was used for inference without quantization, which allowed it to deliver results with full precision. This approach was ideal when computational resources were sufficient and precision was critical. On the other hand, the Llama3-70B 101010meta-llama/Meta-Llama-3-70B-Instruct model, due to hardware limitations, underwent quantization. This process reduced the model’s size and computational needs, making it accessible despite the constraints of our hardware. While quantization might lead to some loss of precision, it was necessary to deploy a model with more parameters, which otherwise would not have been feasible.
@@
Jiang et al. (2023a) In our experiments, we used the Mistral-7B-Instruct-v0.2111111mistralai/Mistral-7B-Instruct-v0.2 and the Mixtral-8x7B-Instruct-v0.1121212mistralai/Mixtral-8x7B-Instruct-v0.1 in their unquantized forms. The Mistral 7B-Instruct v0.2 was selected for its strong instruction-following abilities, maintaining full precision. Similarly, the Mixtral-8x7B-Instruct-v0.1, which integrates eight 7B models, was used unquantized to benefit from its combined performance. Our goal was to explore the effectiveness of smaller models for our riddle tasks, and we preferred these instruction-tuned variations to ensure they were well-suited to our specific requirements, leading to more effective results.
@@ Appendix G Detailed Results
In this section, we present the detailed results for both datasets across all our experimental techniques. Due to the extensive nature of the experiments, the results for each dataset are organized into subtables. We begin with the BrainTeaser dataset, where the metrics are more detailed due to the dataset’s structure, followed by the results for the RiddleSense dataset. The structure is organized by method per model, with results presented in descending order based on score for each method, rather than by the number of examples. This differs from the approach outlined in 5.

### v4

@@ Appendix F Experimental Setup
AI@Meta (2024) In our experiments, we chose two variations of the model: the 8B and 70B versions. The Llama3-8B101010meta-llama/Meta-Llama-3-8B-Instruct model was used for inference without quantization, which allowed it to deliver results with full precision. This approach was ideal when computational resources were sufficient and precision was critical. On the other hand, the Llama3-70B 111111meta-llama/Meta-Llama-3-70B-Instruct model, due to hardware limitations, underwent quantization. This process reduced the model’s size and computational needs, making it accessible despite the constraints of our hardware. While quantization might lead to some loss of precision, it was necessary to deploy a model with more parameters, which otherwise would not have been feasible.
@@
Jiang et al. (2023a) In our experiments, we used the Mistral-7B-Instruct-v0.2121212mistralai/Mistral-7B-Instruct-v0.2 and the Mixtral-8x7B-Instruct-v0.1131313mistralai/Mixtral-8x7B-Instruct-v0.1 in their unquantized forms. The Mistral 7B-Instruct v0.2 was selected for its strong instruction-following abilities, maintaining full precision. Similarly, the Mixtral-8x7B-Instruct-v0.1, which integrates eight 7B models, was used unquantized to benefit from its combined performance. Our goal was to explore the effectiveness of smaller models for our riddle tasks, and we preferred these instruction-tuned variations to ensure they were well-suited to our specific requirements, leading to more effective results.
@@ Appendix G Detailed Results
In this section, we present the detailed results for both datasets across all our experimental techniques. Due to the extensive nature of the experiments, the results for each dataset are organized into subtables. We begin with the BrainTeaser dataset, where the metrics are more detailed due to the dataset’s structure, followed by the results for the RiddleSense dataset. The structure is organized by method per model, with results presented in descending order based on score for each method, rather than by the number of examples. This differs from the approach outlined in Section 5.